OpenStudy (anonymous):

If h(x)=x-5, find a function g(x) such that g(h(x))=x. g(x)=?

OpenStudy (anonymous):

Sounds like an inverse function problem. Is that correct?

OpenStudy (anonymous):

I think so. That is what I am trying right now.

OpenStudy (anonymous):

Original equation is y= x-5. Here, we must solve for x in terms of y. So, x= y + 5. To make it the inverse, switch x and y. y = x+5

OpenStudy (anonymous):

That is what I got the first time but when it graded it it said it was wrong. But I tried again and it said it was right. Thanks!